# MnLargeSymbols’164 MnLargeSymbols’171

### v1

@@ 7.2   Encodings
In the amplitude encoding case, the initial state is prepared by mapping the features onto the computational basis, thus taking advantage of a bigger portion of the Hilbert space. For this reason, one can afford much larger latent spaces (or, equivalently, reduce the demand of qubits) that can be exploited to better identify the different classes in the original dataset. As such, one expects the amplitude encoding to outperform the other methods for a sufficiently large number of features. We show in Fig. 4 the accuracy of the quantum classifier with the amplitude encoding as a function of the dimension of the latent space. For definiteness, we fix the number of qubits to $N=7$ and we show, for comparison, the performances achieved with the other encodings.
@@
Figure 4: Training (left panel) and testing (right panel) accuracy as a function of the number of features in the latent space obtained with the amplitude encoding with $N=7$. The feature reduction has been performed with the autoencoder $AE_{1}$ and the time evolution with the Hamiltonian $H_{1}$. Dashed horizontal lines represent the accuracy achieved with the other encoding schemes.
@@ 7.3   Hamiltonians
Finally, we fix the feature reduction strategy and the encoding method and we investigate the performance of the quantum classifier by comparing the three different Hamiltonians introduced in Sec. 5. The accuracy rapidly grows with increasing the number of qubits and reaches a plateau around $N=10$. The results are not sensitive to the specific Hamiltonian.
@@
This lack of sensitivity can be attributed to the complex nature of the dynamics generated by the three Hamiltonians, which ensures that the system explores a large set of states within the Hilbert space over time. Consequently, the overall performance becomes independent of the specific details of the Hamiltonian employed. This intuition is consistent with the observation that when non-interacting Hamiltonians are considered and the system remains separable throughout its evolution, the accuracy drops significantly. For comparison, we also show the classification accuracy obtained after stripping off the autoencoder and the quantum layer, namely by a purely classical ONN which is fed with the full image information (the 784 pixel values). The classical and quantum classifiers achieve similar results for $N=6$, but the latter clearly outperforms if more qubits are employed.
@@
Figure 5: Training (left panel) and testing (right panel) accuracy as a function of the number of qubits for three different Hamiltonians. The feature reduction and the encoding have been performed with the autoencoder $AE_{1}$ and the dense angle, respectively. The horizontal line represents the accuracy obtained with a ONN fed with all the pixel values of the images.
@@ 8   Conclusions
In this work, we have explored several implementation aspects of quantum machine learning techniques for image classification problems. We went through the full quantum machine learning pipeline, from the preparation of the classical data to the final classification task.
For the sake of concreteness, and to identify the most promising methodologies in a simple environment, we considered the hand-written digits MNIST dataset for a multilabel classification problem. This model can be easily extended to more complex and realistic datasets.
The first step in the workflow consists in the reduction of the features of the original data into a compressed representation. This is an essential step to comply with the limited number of qubits available/simulable with the current technology.
@@
We inspected five different encoding strategies usually considered in the literature. Four of them operate on each separate qubit and another one spans the whole Hilbert space. We found that, among the former, the dense angle and the uniform Bloch sphere encodings provide the best performance and, in contrast to the amplitude encoding, the easiest practical implementation on real quantum circuits. This result is particularly encouraging since it naturally maximizes the trade-off between performance and feasibility.
On the other hand, although amplitude encoding may be difficult to realize on real hardware, for the same number of qubits, it maps more features compared to the singlet-qubit methods and, therefore, can provide better accuracies.
Finally, we studied the time evolution governed by three different Hamiltonians. The first one is characterized by an explicit time dependence, while the other two are time-independent. Of the latter, one is characterized by a system of qubits interacting among of all them, and the other in which interactions are effective only among closest neighbors. Remarkably, the classification performance is effectively the same for the three Hamiltonians, thus suggesting that we could safely opt for the one that is easier to implement without affecting the final gain.
For the final classification task, we used a single-layer fully-connect neural network, in line with the extreme machine learning prescription.
The present analysis could be extended in several directions. For instance, one could define several performance metrics (e.g. classification accuracy, time complexity, energy consumption, etc., and/or combinations of them) to compare the QELM algorithms against different classical models. This could be useful to quantify the quantum advantage. Moreover, one could explore the actual implementability of the models with real hardware.
@@ Acknowledgments
The results have received funds from the program Plan de Doctorados Industriales of the Research and Universities Department of the Catalan Government (2022 DI 011). The research leading to these results has received funding from the Spanish Ministry of Science and Innovation PID2022-136297NB-I00 /AEI/10.13039/501100011033/ FEDER, UE. IFAE is partially funded by the CERCA program of the Generalitat de Catalunya. JS acknowledges the contribution from PRIN (Progetti di Rilevante Interesse Nazionale) TURBIMECS - Turbulence in Mediterranean cyclonic events, grant n. 2022S3RSCT CUP H53D23001630006, CUP Master B53D23007500006.

### v3

@@ 7.2   Encodings
In the amplitude encoding case, the initial state is prepared by mapping the features onto the computational basis, thus taking advantage of a bigger portion of the Hilbert space. For this reason, one can afford much larger latent spaces (or, equivalently, reduce the demand of qubits) that can be exploited to better identify the different classes in the original dataset. As such, one expects the amplitude encoding to outperform the other methods for a sufficiently large number of features and fixed and small number of qubits. We show in Fig.4 the accuracy of the quantum classifier with the amplitude encoding as a function of the dimension of the latent space. For definiteness, we fix the number of qubits to $N=7$ and we show, for comparison, the performances achieved with the other encodings.
@@
Figure 4: Training (left panel) and testing (right panel) accuracy as a function of the number of features in the latent space obtained with the amplitude encoding with $N=7$. The feature reduction has been performed with the autoencoder $AE_{1}$ and the time evolution with the Hamiltonian $H_{1}$. Dashed horizontal lines represent the accuracy achieved with the other encoding schemes with the same number of qubits, $N=7$.
@@ 7.3   Hamiltonians
Finally, we fix the feature reduction strategy and the encoding method and investigate the performance of the quantum classifier by comparing the different Hamiltonians introduced in Sec.5. In detail, we can see in Fig.5 similar and good accuracy performance as function of the number of qubits for the time-dependent Hamiltonian $H_{1}$, the transverse-field Ising Hamiltonians $H_{2}$ and $H_{3}$, the Heisenberg XXZ model $H_{4}$, the integrable and non-ergodic XX $H_{5}$ and the Hamiltonian $H_{6}$ with parameters in the non-ergodic and transition region. The accuracy rapidly grows with increasing the number of qubits and reaches a plateau around $N=10$.
This lack of sensitivity can be attributed to the complex nature of the dynamics generated by these Hamiltonians, it spreads the information from the local states (where the encoding occurs) to the computational basis, combining the different pieces of information encoded in different qubits. When the probability of the outcomes is measured, the recombined information is extracted, allowing the device to learn from the process.
This, however, does not occur, for instance, when Hamiltonians that induce localization are used. We show in Fig. 6 the performance accuracy as function of the number of qubits for the Hamiltonians $H_{6}$ in localizations areas (region I and III of Ref. [53]) and for transverse-field Ising Hamiltonian $H_{2}$ when $J_{2}=0$, that is where there are no interactions among the spins. In the Fig. 6, we have included $H_{2}$ again, as a reference for the better performance of the Hamiltonians, and here we can notice how $H_{6}$ and $H_{2}(J_{2}=0)$ have worse performance.
@@
For comparison, in Fig. 5 and Fig. 6, we also show the classification accuracy obtained after stripping off the autoencoder and the quantum layer; namely, by a purely classical ONN which is fed with the full image information (the 784 pixel values). The classical and quantum classifiers achieve similar results for $N=6$, but the latter clearly outperforms if more qubits are employed.
Additionally, in Fig. 5 we present a further classical comparison with an ONN directly fed with the features reduced by the autoencoder $AE_{1}$ (with a latent space dimension given by $2N$). Also in this case, after $N=6$, the quantum classifier outperforms the classical one.
The accuracy obtained with the ONN using the raw 784 data is in general greater than, or comparable with, the one achieved with the compressed features (coral line), the latter slightly exceeding the former when $N=11$ for the train set and $N=8$ for the test set. This suggests that, for a straightforward comparison with the quantum model, the ONN with the raw data represents a simple and viable choice, at least for the range of $N$ analyzed in this work.
Furthermore, it is important to emphasize that the task of the autoencoder is not the classification itself, but rather the reduction of the amount of available information. While in some cases this may help to remove the noise, in general it leads to a loss of data. The use of the AE in the quantum model is not a strategic advantage over the classical model in the classification task, but rather a necessity due to the current limitations of quantum hardware, which can only handle a limited number of qubits.
To further investigate the role of the autoencoder and, more generally, of a compressed representation, one could devise in principle a quantum algorithm to which the raw and uncompressed data is directly provided as input. While this cannot be achieved using the QELM architectures developed in our work, due to the aforementioned computational limitations, the importance of the compressed representations, as compared to the uncompressed one, can still be highlighted. Indeed, we are able to encode the full data using the amplitude encoding, provided that $N\geq 10$ (the other encodings would require at least $N=392$ qubits to store the entire image). In this case, the accuracy on the test sets, using $H_{1}$ as an example, is found to be $\simeq 0.968$ for $N=10$, $0.975$ for $N=11$ and $0.976$ for $N=12$. These accuracies are smaller than those obtained, using the same number of qubits ($N=10,11,12$), with the dense angle encoding, for which the compressed representation of the dataset is strictly necessary (see for instance Fig. 5(b)). Besides the actual performance of the algorithms, we want to emphasize once more, that the use of the compressed representation is way more efficient as it requires less resources while preserving as much information as possible of the original data.
Finally, Fig. 5(a) and Fig. 5(b) clearly show that the larger accuracies obtained with the quantum algorithm are not due to the dimensionality reduction, but to the presence of the quantum layer.
@@
Figure 5: Training (left panel) and testing (right panel) accuracy as a function of the number of qubits for six different Hamiltonians. The feature reduction and the encoding have been performed with the autoencoder $AE_{1}$ and the dense angle encoding, respectively. The horizontal line represents the accuracy obtained with a ONN fed with all the pixel values of the images. Instead, the coral line represents a ONN fed directly with the $2N$ features extracted by the $AE_{1}$.
Figure 6: Training (left panel) and testing (right panel) accuracy as a function of the number of qubits for other different Hamiltonians. The feature reduction and the encoding have been performed with the autoencoder $AE_{1}$ and the dense angle encoding, respectively. The horizontal line represents the accuracy obtained with a ONN fed with all the pixel values of the images.
@@ 8   Conclusions
In this work, we have explored several implementational aspects of quantum machine learning techniques for image classification problems. We went through the full quantum machine learning pipeline, from the preparation of the classical data to the final classification task.
To demonstrate our methodology in a concrete manner and identify the most promising approaches in a controlled setting, we employed the MNIST handwritten digits dataset for a multilabel classification problem. Moreover, our experiments with the Fashion-MNIST dataset confirm that this model can be readily extended to more complex and realistic scenarios.
The first step in the workflow consists of the reduction of the features of the original data into a compressed representation. This is an essential step to comply with the limited number of qubits available/simulable with the current technology.
@@
We inspected five different encoding strategies typically considered in the literature. Four of them operate on each qubit separately, while another one spans the whole Hilbert space. We found that, among the former, the dense angle and the uniform Bloch sphere encodings provide the best performance and, in contrast to the amplitude encoding, enjoy the easiest practical implementation on real quantum circuits. This result is particularly encouraging, as it naturally maximizes the trade-off between performance and feasibility.
On the other hand, although amplitude encoding may be difficult to realize on real hardware, it maps more features compared to the single-qubit methods and, therefore, can provide better accuracies if the number of qubits is small.
Finally, we studied the time evolution governed by six different Hamiltonians. The first one is characterized by an explicit time dependence, while the other five are time-independent. Of the latter, two are characterized by a system of qubits with all-to-all interactions, while others involve interactions effective only among closest neighbors. Moreover, among them, there are Hamiltonians we have tested to explore the role of ergodicity, integrability, and spin-spin interaction in the performance of the QELM.
We found that integrability and ergodicity do not play a distinguishing role in themselves; rather, it seems that a key factor is the spread of information from the initial local states to the final measurement states in the computational bases, passing through the eigenstates of the Hamiltonian. This does not work well, for example, in the case of localization.
Remarkably, the classification performance is effectively the same for the different Hamiltonians, thus suggesting that we could safely opt for the one that is easier to implement without affecting the final gain.
For the final classification task, we used a single-layer fully-connected neural network, in line with the extreme learning machine prescription.
The present analysis could be extended in several directions. For instance, one could define several performance metrics (e.g., classification accuracy, time complexity, energy consumption, etc., and/or combinations of them) to compare the QELM algorithms against different classical models. This could be useful to quantify the quantum advantage. Moreover, one could explore the actual implementability of the models with real hardware.
@@ Acknowledgments
The results have received funds from the program Plan de Doctorados Industriales of the Research and Universities Department of the Catalan Government (2022 DI 011). The research leading to these results has received funding from the Spanish Ministry of Science and Innovation PID2022-136297NB-I00 /AEI/10.13039/501100011033/ FEDER, UE. IFAE is partially funded by the CERCA program of the Generalitat de Catalunya. JS acknowledges the contribution from PRIN (Progetti di Rilevante Interesse Nazionale) TURBIMECS - Turbulence in Mediterranean cyclonic events, grant n. 2022S3RSCT CUP H53D23001630006, CUP Master B53D23007500006. FP and NL aknowledge support by the PNRR MUR Project N. PE0000023-NQSTI.
@@ Appendix A Appendix
Appendix B Appendix
To assess the robustness of our QELM algorithm, we also tested it on a more complex dataset, Fashion-MNIST. This dataset contains 70000 28x28 grayscale images of 10 different fashion categories, including clothing items such as t-shirts, shoes, and dresses.
We reduced the images using either the autoencoder $AE_{1}$ or PCA and fed them into our algorithm for image classification. The conclusions that can be drawn from the results in Fig. 8 are consistent with those found for the MNIST dataset and the gap between the accuracies of $AE_{1}$ (green line) and PCA (red line) persists in the range of qubits studied in this work.
Figure 8: Training (solid line) and testing (dashed line) accuracy as a function of the number of qubits for different feature reduction schemes for the Fashion-MNIST dataset: PCA (red) and autoencoder $AE_{1}$ (green). The time evolution and the encoding have been performed with the Hamiltonian $H_{1}$ and the dense angle encoding, respectively.